# MyFramelinecolor=blue, outerlinewidth=0.7pt, roundcorner=12pt, innertopmargin=.5nerbottommargin=.5nerrightmargin=1pt, innerleftmargin=1pt, backgroundcolor=white

### v1

@@ 1.1. Intrinsic enumerative mirror symmetry
with $\tau,\sigma$ $A$-model symplectic structure parameters and $s,t$ $B$-model complex structure parameters.
@@ 1.2. Outline and main result
$A$-model and Takahashi’s mirror conjecture (§2)
@@
As a corollary we deduce Takahashi’s enumerative mirror conjecture in §4.5. We emphasize that the proof of Theorem 1.1 is by direct and rather trivial computation. Moreover, the matching of symplectic and Kähler parameters is already built into intrinsic mirror symmetry.
@@ 2.1. $A$-model
Let $\overline{M}_{0,(3d)}(\mathbb{P}^{2},E,d)$ be the moduli space of genus 0 degree $d$ basic stable log maps to $(\mathbb{P}^{2},E)$ meeting $E$ in one unspecified point of maximal tangency $3d$ [Ch, AC, GS13]. It is of virtual dimension 00 and admits a virtual fundamental class $[\overline{M}_{0,(3d)}(\mathbb{P}^{2},E,d)]^{\rm vir}\in{\rm H}_{0}\left(\overline{M}_{0,(3d)}(\mathbb{P}^{2},E,d),\mathbb{Q}\right)$. We obtain the degree $d$ maximal tangency genus 0 log Gromov–Witten invariants
@@ 2.2. Takahashi’s mirror family
We review the setup from [T01]. Takahashi constructs the mirror family from toric duality principles. The fan of $\mathbb{P}^{2}$ is generated by $(1,0),\,(0,1),\,(-1,-1)\in\mathbb{Z}^{2}\otimes_{\mathbb{Z}}\mathbb{R}$. Let $\Delta^{\circ}$ be the reflexive polygon generated by $(1,0),\,(0,1),\,(-1,-1)$. Then the associated toric variety is $\mathbb{P}_{\Delta^{\circ}}\cong\mathbb{P}^{2}\big{/}\mu_{3}$, where $\lambda\in\mu_{3}=\{1,e^{\pm 2\pi i/3}\}$ acts on homogeneous coordinates by
@@ 2.3. Enumerative Mirror Conjecture
Conjecture 1.10 from [T01], as noted in Remark 1.11 of the same source, asserts that, using the canonical coordinate $q$ defined by
@@ 3. Intrinsic mirror family
In the Fano case, $R$ is the completion of a finitely generated $\mathbb{C}\llbracket P\rrbracket$-algebra, so one naturally obtains an algebraic family
@@ 3.1. Starting point: A maximal degeneration of $(\mathbb{P}^{2},E)$
To obtain a family of two-dimensional varieties as a mirror, we rather need to apply the mirror construction to a simple normal crossing degeneration $g:(\mathcal{Y},\mathcal{D})\to C$ over a curve $C$ with general fiber $(\mathbb{P}^{2},E)$ and $\mathcal{D}$ with a zero-dimensional stratum. Thus, $\mathcal{D}$ is a simple normal crossings divisor in a smooth $\mathcal{Y}$ surjecting to the smooth curve $C$ and with $\mathcal{Y}_{0}=g^{-1}(0)\subseteq\mathcal{D}$ for a special closed point $0\in C$. The remaining component of $\mathcal{D}$ surjects onto $C$ and has general fiber an elliptic curve. To construct $\mathcal{Y}$ we consider the family $\overline{\mathcal{Y}}$ of hypersurfaces
$XYZ-s(U+f)=0$
@@
This family is easy to understand by viewing $\mathbb{P}(1,1,1,3)$ as the toric variety with momentum polytope with vertices $(-1,-1,0)$, $(2,-1,0)$, $(-1,2,0)$, $(0,0,1)$. This is a tetrahedron with one facet $\sigma$ a standard $2$-simplex scaled by $3$, and the vertex $v$ not in this face has integral distance $1$ from this face. The monomial $U$ corresponds to $v$, and $X^{3},Y^{3},Z^{3},XYZ$ to the vertices and only interior integral point in $\sigma$. For $s\neq 0$ we can eliminate $U$ to obtain $\mathbb{P}^{2}$ with the standard homogenous coordinates $X,Y,Z$. The divisor $U=0$ leads to the family $XYZ-sf(X,Y,Z)=0$ of smooth cubic curves.
@@
minus the four singular members, each a union of three lines, at $\lambda/\mu\in A=\{0,\omega,\omega^{2},1\}$ for $\omega$ a primitive third root of unity. The base locus of the Hesse pencil is the set of nine flex points of each member, so the flex points do not depend on $[\lambda,\mu]$. The automorphism group $\Gamma$ of the Hesse configuration induces an action on the Hesse pencil. The induced action on the parameter space $\mathbb{P}^{1}$ factors over a faithful action of the alternating group $A_{4}$ permuting the critical values. See [AD] for a comprehensive discussion. This shows
@@ 3.3. The tropicalization of $(\mathcal{Y},\mathcal{D})$
Applying the tropicalization construction to $(\mathcal{Y},\mathcal{D})$ gives three standard simplicial cones $\sigma_{1},\sigma_{2},\sigma_{3}=\mathbb{R}_{\geq 0}^{3}$ for the three normal crossing points on $\mathcal{Y}_{0}$ depicted at the boundary of the large triangles in Figure 3.1, while a local toric computation at the remaining zero-dimensional stratum of $\mathcal{Y}_{0}$ gives
@@
$\overline{\sigma}_{3}$$\overline{\sigma}_{0}$$\overline{\sigma}_{1}$$\overline{\sigma}_{2}$$e_{1}$$e_{3}$$e_{2}$
@@ 3.4. The integral affine manifold $B$ and asymptotic geometry
$\overline{\sigma}_{3}$$\overline{\sigma}_{2}$$\overline{\sigma}_{1}$
@@
The vertices are at positions $(-3,-1),(0,0),(0,1),(-3,2)$. The diagram could be periodically continued vertically to yield a chart on the universal cover of $B\setminus\overline{\sigma}_{0}$ via powers of the affine coordinate transformation $x\mapsto A\cdot x+b$ with
@@ 3.5. Wall structures and intrinsic mirror families
In [GS19], the intrinsic mirror ring of $(\mathcal{Y},\mathcal{D})$ is defined directly as the free $\mathbb{C}\llbracket P\rrbracket$-module
@@
The structure coefficients $N_{pqr}^{A}\in\mathbb{Q}$ are logarithmic Gromov-Witten counts of genus zero stable log maps with curve class $A$ and contact orders $p,q$ and $-r$ with $\mathcal{D}$. The negative contact order $-r$ is implemented by the theory of punctured logarithmic maps [ACGS]. If $\tau$ is the minimal cone containing $r$ then the irreducible component containing the marked point of the domain of a punctured stable map with contact order $-r$ contributing to $N_{pqr}^{A}$ maps into the closed stratum $\mathcal{Y}_{\tau}$. The main result of [GS19] is that (3.6) defines a ring structure on $R$, and in particular is associative. For the grading of $R$ one takes $\deg\vartheta_{p}$ as the integral length of $g(p)$ for $g:|\Sigma(\mathcal{Y})|\to\mathbb{R}_{\geq 0}$ as in (3.2), the contact order with the central fiber $\mathcal{Y}_{0}$.
Better control of intrinsic mirrors, at the expense of slightly more restrictive assumptions, which hold in all cases relevant to mirror symmetry, is provided by the alternative construction of $R$ in [GS22]. This approach employs the previously developed machinery of wall structures and generalized theta functions [GS11, GPS, GHK, GHS]. A wall $\mathfrak{p}$ in this setup is a codimension one rational polyhedral subset of $|\Sigma(\mathcal{Y})|$ contained in one cone $\sigma$ of $\Sigma(\mathcal{Y})$ along with an element $f_{\mathfrak{p}}$ of the Laurent polynomial ring $\mathbb{C}\llbracket P\rrbracket[\mathfrak{p}_{\mathbb{Z}}^{\operatorname{gp}}]$ with exponents integral tangent vector fields on $\mathfrak{p}$ and coefficients in $\mathbb{C}\llbracket P\rrbracket$, the completed ring of curve classes. The construction requires $f_{\mathfrak{p}}$ to be homogeneous of degree 00, so all exponents are contracted by $g|_{\sigma}$, and $f_{\mathfrak{p}}\equiv 1$ modulo the maximal ideal $\mathfrak{m}=(P\setminus\{0\})$ of $\mathbb{C}\llbracket P\rrbracket$.
@@
see [GS22, (3.9)–(3.11)]. Here $-u_{\mathfrak{t}}\in\sigma_{\mathbb{Z}}\subseteq\Sigma(\mathcal{Y})(\mathbb{Z})$ is the contact order at the puncture, $\sigma$ is the smallest cone containing $\mathfrak{p}$, and $k_{\tau}\in\mathbb{N}\setminus\{0\}$ is a tropical multiplicity. The set of all such walls defines the canonical wall structure of [GS22]. Figure 3.4 shows the canonical wall structure of $(\mathcal{Y},\mathcal{D})$ in the asymptotic chart of Figure 3.3 up to curves of degree $3$.
@@
Here we use that the closed stratum $\mathcal{Y}_{\rho}$ is one-dimensional, hence has an associated class $[\mathcal{Y}_{\rho}]\in P$. If $\sigma,\sigma^{\prime}$ are the maximal cones in $\Sigma(\mathcal{Y})$ containing $\rho$ let $\xi\in(\sigma-\rho)_{\mathbb{Z}}$ be contracted by $g$ and generate a complementary summand to $\rho_{\mathbb{Z}}^{\operatorname{gp}}$ in $\sigma_{\mathbb{Z}}^{\operatorname{gp}}$. Mapping $u$ to $z^{\xi}$ and $v$ to $f_{\mathfrak{b}}t^{[\mathcal{Y}_{\rho}]}z^{-\xi}$ induces an isomorphism of the localization
@@
To extend this flat deformation to all of $X_{0}$ requires in addition consistency of the wall structure in codimension [GHS, Def. 3.2.1]. This condition amounts to saying that the restriction of $X_{I}^{\circ}$ to an affine subset $U\subseteq X_{0}$ has enough regular functions to induce a flat deformation $U_{I}$ of $U$. One then obtains a flat deformation of a scheme that is projective over an affine scheme with central fiber $X_{0}$, and a distinguished $S_{I}$-module basis $\vartheta_{p}^{I}$ of its homogeneous coordinate ring $R_{I}$ for an ample line bundle coming with the construction. There is one generator $\vartheta^{I}_{p}\in R_{I}$ for each contact order $p\in\Sigma(\mathcal{Y})(\mathbb{Z})$, and these are compatible with changing $I$.
Another key result of [GS22] is that $R_{I}$ agrees with the reduction modulo $I$ of the intrinsic mirror ring $R$, for all $I$:
@@
Analyzing curves of low degree and using the uniqueness of consistent wall structures [GS11], one can show that up to a trivial change [^3] The only difference is that the singularities of the affine structure on $B$ are moved from the interior of the bounded edges to the vertices, see [Gr20]. the canonical wall structure agrees with the algorithmically constructed wall structure from [GS11, CPS], see [Gr22a]. For the sequel, we only need the following statement.
@@
The theta function $\vartheta_{p}$ is defined as an element of the ring $R_{\mathfrak{u}}=S_{I}[\sigma_{\mathbb{Z}}^{\operatorname{gp}}]$ for a chamber $\mathfrak{u}$ in a maximal cone $\sigma$ in $\Sigma(\mathcal{Y})$ by a sum over so-called broken lines with initial direction $-p$ and ending at a general, specified point $x\in\operatorname{Int}\mathfrak{u}$. A broken line $\beta$ is a piecewise straight path in $|\Sigma(\mathcal{Y})|$, contained in the complement of the codimension two skeleton and carrying a monomial $c\cdot z^{m}$ on each straight line segment, with $m$ tangent to $\beta$ and pointing backwards. So $m=-p$ on the unique unbounded line segment of $\beta$. When $\beta$ meets a wall $\mathfrak{p}$, one extends $\beta$ across the wall with one of the summands in the expansion of $\theta_{\mathfrak{p}}(cz^{m})$ into a sum of Laurent monomials. We refer to [GHS, §3.1] for the precise definition. In particular, the exponent $m$ carried by $\beta$, hence the direction of $\beta$, can only change at a wall by adding a multiple of one of the exponents appearing in $f_{\mathfrak{p}}$. The contribution of each broken line to $\vartheta_{p}$ for the local expression in $S_{I}[\sigma_{\mathbb{Z}}^{\operatorname{gp}}]$ is the monomial $cz^{m}$ at its endpoint $x\in\operatorname{Int}\sigma$.
From the outward pointing nature of the wall structure (Lemma 3.2), it is then easy to see that the only broken lines contributing to $\vartheta_{0},\ldots,\vartheta_{3}$ at a point close to the barycentric ray of $\sigma_{0}$ are straight (no bend), while those contributing to $\vartheta_{4}$ bend at most once when crossing from an unbounded chamber into $\sigma_{0}$, by a primitive integral tangent vector of the crossed facet of $\sigma_{0}$. The first relation in (3.13) therefore holds to all orders, while the relation involving $\vartheta_{4}$ for the chosen symmetric resolution $\mathcal{Y}\to\overline{\mathcal{Y}}$ turns out to give the Hori-Vafa mirror superpotential [CPS, Cor. 7.9].
@@
The intrinsic mirror of $(\mathcal{Y},\mathcal{D})$, after the base change $\mathbb{C}\llbracket P\rrbracket\to\mathbb{C}\llbracket t\rrbracket$ given by $[C]\mapsto[C]\cdot D_{4}$, equals
@@
Here $X=\vartheta_{1}$, $Y=\vartheta_{2}$, $Z=\vartheta_{3}$, $U=\vartheta_{0}$ are of degree $1$ and $W=\vartheta_{4}$ is of degree 00.
@@
fulfilling a compatibility condition [GHS, (4.8)]. Here $\operatorname{PL}(B)$ is the monoid of continuous maps $B\to\mathbb{R}$ that are piecewise affine with integral slopes on each cell of the polyhedral decomposition $\mathscr{P}$ of $B$. Explicitly, denoting by $\Sigma(\mathcal{Y})[1]$ the set of rays of $\Sigma(\mathcal{Y})$, the dual $\operatorname{PL}(B)^{*}$ is the quotient of $\mathbb{Z}^{\Sigma(\mathcal{Y})[1]}$ by the subspace of tuples pairing trivially with all piecewise linear functions. In our case of $Q=P=\mathbb{N}$ we take $\delta_{Q}=\operatorname{id}$ and $\delta_{B}$ the evaluation at the PL-function $\delta$ from Lemmma 3.2 that vanishes on all bounded cells of the polyhedral decomposition $\mathscr{P}$ of $B$ and has slope $1$ on the unbounded rays.
It remains to check that the wall functions from (3.7) are homogeneous of degree 00 for the degree of monomials defined from $\delta_{Q}$ and $\delta_{B}$. Each wall function is a function in a monomial $t^{A}z^{-u_{\tau}}$ for a curve class $A$ and outgoing contact order $u_{\tau}$ for a type $\tau$ of a $1$-punctured map. In the notation of [GHS], this monomial has exponent $m=(\overline{m},m_{Q})$ with $\overline{m}=-u_{\tau}$, $m_{Q}=A\in P=\mathbb{N}$, and degree
@@
We have $\delta_{Q}(m_{Q})=A\in\mathbb{N}$, the intersection number $[C]\cdot\mathcal{E}$ for the class of a curve $C$ contributing to the wall. In the second summand, $\nabla_{\overline{m}}\in\operatorname{PL}(B)^{*}$ is the directional derivative along $\overline{m}=-u_{\tau}$, so $\delta_{B}(\nabla_{\overline{m}})=\nabla_{\overline{m}}(\varphi)$. The balancing condition [ACGS, Cor. 2.30] applied with $s$ the section of ¯$\overline{\mathcal{M}}_{X}^{\operatorname{gp}}$ defined by $\mathcal{E}$ then shows

### v2

# Intrinsic enumerative mirror symmetry: Takahashi’s log mirror symmetry for $(\mathbb{P}^{2},E)$ revisited

@@ 1.1. Intrinsic enumerative mirror symmetry
with $\tau,\sigma$ AA-model symplectic structure parameters and $s,t$ $B$-model complex structure parameters.
@@ 1.2. Outline and main result
AA-model and Takahashi’s mirror conjecture (§2)
@@
As a corollary, we deduce Takahashi’s enumerative mirror conjecture in §4.5. We emphasize that the proof of Theorem 1.1 is by direct and rather trivial computation. Moreover, the matching of symplectic and Kähler parameters is already built into intrinsic mirror symmetry.
@@ 2.1. $A$-model
Let $\overline{M}_{0,(3d)}(\mathbb{P}^{2},E,d)$ be the moduli space of genus 0 degree $d$ basic stable log maps to $(\mathbb{P}^{2},E)$ meeting $E$ in one unspecified point of maximal tangency $3d$ [Ch, AC, GS13]. It is of virtual dimension 0 and admits a virtual fundamental class $\overline{M}_{0,(3d)}(\mathbb{P}^{2},E,d)]^{\rm vir}\in{\rm H}_{0}\left(\overline{M}_{0,(3d)}(\mathbb{P}^{2},E,d),\mathbb{Q}\right)$. We obtain the degree $d$ maximal tangency genus 0 log Gromov–Witten invariants
@@ 2.2. Takahashi’s mirror family
We review the setup from [T01]. Takahashi constructs the mirror family from toric duality principles. The fan of $\mathbb{P}^{2}$ is generated by $(1,0),\,(0,1),\,(-1,-1)\in\mathbb{Z}^{2}\otimes_{\mathbb{Z}}\mathbb{R}$. Let $\Delta^{\circ}$ be the reflexive polygon with vertices (1,0),(0,1),(1,1)(1,0),\,(0,1),\,(-1,-1). Then the associated toric variety is $\mathbb{P}_{\Delta^{\circ}}\cong\mathbb{P}^{2}\big/\mu_{3}$, where $\lambda\in\mu_{3}=\{1,e^{\pm 2\pi i/3}\}$ acts on homogeneous coordinates by
@@ 2.3. Enumerative Mirror Conjecture
Conjecture 1.10 from [T01] together with Remark 1.11 of the same source, asserts that, using the canonical coordinate $q$ defined by
@@ 3. Intrinsic mirror family
If $D$ supports a nef divisor, $R$ is the completion of a finitely generated $\mathbb{C}\llbracket P\rrbracket$-algebra [GS19, Constr. 1.24], so one naturally obtains an algebraic family
@@
If $D$ even supports an ample divisor, $R$ is of finite type over $\mathbb{C}[P]$ [GS19, Rem. 1.26]. Thus in this case, the mirror family is given by polynomial equations.
@@ 3.1. Starting point: A maximal degeneration of $(\mathbb{P}^{2},E)$
To obtain a family of two-dimensional varieties as a mirror, we rather need to apply the mirror construction to a simple normal crossings degeneration $g:(\mathcal{Y},\mathcal{D})\to C$ over a curve $C$ with general fiber $(\mathbb{P}^{2},E)$ and $\mathcal{D}$ with a zero-dimensional stratum. Thus, $\mathcal{D}$ is a simple normal crossings divisor in a smooth $\mathcal{Y}$ surjecting to the smooth curve $C$ and with $\mathcal{Y}_{0}=g^{-1}(0)\subseteq\mathcal{D}$ for a special closed point $0\in C$. The remaining component of $\mathcal{D}$ surjects onto $C$ and has general fiber an elliptic curve. To construct $\mathcal{Y}$, we consider the family $\overline{\mathcal{Y}}$ of hypersurfaces
XYZs(U+f)=0XYZ-s(U+f)=0
@@
This family is easy to understand by viewing $\mathbb{P}(1,1,1,3)$ as the toric variety with momentum polytope with vertices (1,1,0)(-1,-1,0), (2,1,0)(2,-1,0), (1,2,0)(-1,2,0), $(0,0,1)$. This is a tetrahedron with one facet $\sigma$ a standard $2$-simplex scaled by $3$, and the vertex $v$ not in this face has integral distance $1$ from this face. The monomial $U$ corresponds to $v$, and $X^{3},Y^{3},Z^{3},XYZ$ to the vertices and only interior integral point in $\sigma$. For $s\neq 0$ we can eliminate $U$ to obtain $\mathbb{P}^{2}$ with the standard homogenous coordinates $X,Y,Z$. The divisor $U=0$ leads to the family XYZsf(X,Y,Z)=0XYZ-sf(X,Y,Z)=0 of smooth cubic curves.
@@
minus the four singular members, each a union of three lines, at $\lambda/\mu\in A=\{0,\omega,\omega^{2},1\}$ for $\omega$ a primitive third root of unity. The base locus of the Hesse pencil is the set of nine flex points of each member, so the flex points do not depend on $\lambda,\mu]$. The automorphism group $\Gamma$ of the Hesse configuration induces an action on the Hesse pencil. The induced action on the parameter space $\mathbb{P}^{1}$ factors over a faithful action of the alternating group $A_{4}$ permuting the critical values. See [AD] for a comprehensive discussion. This shows
@@ 3.3. The tropicalization of $(\mathcal{Y},\mathcal{D})$
Applying the tropicalization construction to $(\mathcal{Y},\mathcal{D})$ gives three standard simplicial cones $\sigma_{1},\sigma_{2},\sigma_{3}=\mathbb{R}_{\geq 0}^{3}$ for the three normal crossings points on $\mathcal{Y}_{0}$ depicted as the vertices at the boundary of the large triangles in Figure 3.1, while a local toric computation at the remaining zero-dimensional stratum of $\mathcal{Y}_{0}$ gives
@@
[…]
@@ 3.4. The integral affine manifold $B$ and asymptotic geometry
[…]
@@
The vertices are at positions (3,1),(0,0),(0,1),(3,2)(-3,-1),(0,0),(0,1),(-3,2). The diagram could be periodically continued vertically to yield a chart on the universal cover of $B\setminus\overline{\sigma}_{0}$ via powers of the affine coordinate transformation $x\mapsto A\cdot x+b$ with
@@ 3.5. Wall structures and intrinsic mirror families
To simplify the following discussion of the intrinsic mirror construction, we assume $D$ supports a nef divisor. In [GS19], the intrinsic mirror ring of $(\mathcal{Y},\mathcal{D})$ is then defined directly as the free $\mathbb{C}\llbracket P\rrbracket$-module
@@
The structure coefficients $N_{pqr}^{A}\in\mathbb{Q}$ are logarithmic Gromov-Witten counts of genus zero stable log maps with curve class $A$ and contact orders $p,q$ and r-r with $\mathcal{D}$. The negative contact order r-r is implemented by the theory of punctured logarithmic maps [ACGS]. If $\tau$ is the minimal cone containing $r$ then the irreducible component containing the marked point of the domain of a punctured stable map with contact order r-r contributing to $N_{pqr}^{A}$ maps into the closed stratum $\mathcal{Y}_{\tau}$. The main result of [GS19] is that (3.6) defines a ring structure on $R$, and in particular is associative. For the grading of $R$ one takes $\deg\vartheta_{p}$ as the integral length of $g(p)$ for $g:|\Sigma(\mathcal{Y})|\to\mathbb{R}_{\geq 0}$ as in (3.2), the contact order with the central fiber $\mathcal{Y}_{0}$.
Better control of intrinsic mirrors, at the expense of slightly more restrictive assumptions, which hold in all cases relevant to mirror symmetry, is provided by the alternative construction of $R$ in [GS22]. This approach employs the previously developed machinery of wall structures and generalized theta functions [GS11, GPS, GHK, GHS]. A wall $\mathfrak{p}$ in this setup is a codimension one rational polyhedral subset of $|\Sigma(\mathcal{Y})|$ contained in one cone $\sigma$ of $\Sigma(\mathcal{Y})$ along with an element $f_{\mathfrak{p}}$ of the Laurent polynomial ring $\mathbb{C}\llbracket P\rrbracket[\mathfrak{p}_{\mathbb{Z}}^{\operatorname{gp}}]$ with exponents integral tangent vector fields on $\mathfrak{p}$ and coefficients in $\mathbb{C}\llbracket P\rrbracket$, the completed ring of curve classes. The construction requires $f_{\mathfrak{p}}$ to be homogeneous of degree 0, so all exponents are contracted by $g|_{\sigma}$, and $f_{\mathfrak{p}}\equiv 1$ modulo the maximal ideal $\mathfrak{m}=(P\setminus\{0\})$ of $\mathbb{C}\llbracket P\rrbracket$.
@@
see [GS22, (3.9)–(3.11)]. Here $u_{\mathfrak{t}}\in\sigma_{\mathbb{Z}}\subseteq\Sigma(\mathcal{Y})(\mathbb{Z})$ is the contact order at the puncture, $\sigma$ is the smallest cone containing $\mathfrak{p}$, and $k_{\tau}\in\mathbb{N}\setminus\{0\}$ is a tropical multiplicity. The set of all such walls defines the canonical wall structure of [GS22]. Figure 3.4 shows the canonical wall structure of $(\mathcal{Y},\mathcal{D})$ in the asymptotic chart of Figure 3.3 up to curves of degree $3$.
@@
Here we use that the closed stratum $\mathcal{Y}_{\rho}$ is one-dimensional, hence has an associated class $\mathcal{Y}_{\rho}]\in P$. If $\sigma,\sigma^{\prime}$ are the maximal cones in $\Sigma(\mathcal{Y})$ containing $\rho$ let $\xi\in(\sigma-\rho)_{\mathbb{Z}}$ be contracted by $g$ and generate a complementary summand to $\rho_{\mathbb{Z}}^{\operatorname{gp}}$ in $\sigma_{\mathbb{Z}}^{\operatorname{gp}}$. Mapping $u$ to $z^{\xi}$ and $v$ to $f_{\mathfrak{b}}t^{[\mathcal{Y}_{\rho}]}z^{-\xi}$ induces an isomorphism of the localization
@@
To extend this flat deformation to all of $X_{0}$ requires in addition consistency of the wall structure in codimension two [GHS, Def. 3.2.1]. This condition amounts to saying that the restriction of $X_{I}^{\circ}$ to an affine subset $U\subseteq X_{0}$ has enough regular functions to induce a flat deformation $U_{I}$ of $U$. One then obtains a flat deformation of a scheme that is projective over an affine scheme with central fiber $X_{0}$, and a distinguished $S_{I}$-module basis $\vartheta_{p}^{I}$ of its homogeneous coordinate ring $R_{I}$ for an ample line bundle coming with the construction. There is one generator $\vartheta^{I}_{p}\in R_{I}$ for each contact order $p\in\Sigma(\mathcal{Y})(\mathbb{Z})$, and these are compatible with changing $I$.
Another key result of [GS22] is that $R_{I}$ agrees with the reduction modulo $I$ of the intrinsic mirror ring $R$, for all II:
@@
Analyzing curves of low degree and using the uniqueness of consistent wall structures [GS11, Prop. 4.1], one can show that up to a trivial change [^3] The only difference is that the singularities of the affine structure on $B$ are moved from the interior of the bounded edges to the vertices, see [Gr20]. the canonical wall structure agrees with the algorithmically constructed wall structure from [GS11, CPS], see [Gr22a]. For the sequel, we only need the following statement.
@@
The theta function $\vartheta_{p}$ is defined as an element of the ring $R_{\mathfrak{u}}=S_{I}[\sigma_{\mathbb{Z}}^{\operatorname{gp}}]$ for a chamber $\mathfrak{u}$ in a maximal cone $\sigma$ in $\Sigma(\mathcal{Y})$ by a sum over so-called broken lines with initial direction p-p and ending at a general, specified point $x\in\operatorname{Int}\mathfrak{u}$. A broken line $\beta$ is a piecewise straight path in $|\Sigma(\mathcal{Y})|$, contained in the complement of the codimension two skeleton and carrying a monomial $c\cdot z^{m}$ on each straight line segment, with $m$ tangent to $\beta$ and pointing backwards. So $m=p$ on the unique unbounded line segment of $\beta$. When $\beta$ meets a wall $\mathfrak{p}$, one extends $\beta$ across the wall with one of the summands in the expansion of $\theta_{\mathfrak{p}}(cz^{m})$ into a sum of Laurent monomials. We refer to [GHS, §3.1] for the precise definition. In particular, the exponent $m$ carried by $\beta$, hence the direction of $\beta$, can only change at a wall by adding a multiple of one of the exponents appearing in $f_{\mathfrak{p}}$. The contribution of each broken line to $\vartheta_{p}$ for the local expression in $S_{I}[\sigma_{\mathbb{Z}}^{\operatorname{gp}}]$ is the monomial $cz^{m}$ at its endpoint $x\in\operatorname{Int}\sigma$.
From the outward pointing nature of our wall structure (Lemma 3.2), it is then easy to see that the only broken lines contributing to $\vartheta_{0},\ldots,\vartheta_{3}$ at a point close to the barycentric ray of $\sigma_{0}$ are straight (no bend), while those contributing to $\vartheta_{4}$ bend at most once when crossing from an unbounded chamber into $\sigma_{0}$, by a primitive integral tangent vector of the crossed facet of $\sigma_{0}$. The first relation in (3.13) therefore holds to all orders, while the relation involving $\vartheta_{4}$ for the chosen symmetric resolution $\mathcal{Y}\to\overline{\mathcal{Y}}$ turns out to give the Hori-Vafa mirror superpotential [CPS, Cor. 7.9].
@@
The intrinsic mirror of $(\mathcal{Y},\mathcal{D})$, after the base change $\mathbb{C}\llbracket P\rrbracket\to\mathbb{C}\llbracket t\rrbracket$ given by $C]\mapsto[C]\cdot D_{4}$, equals
@@
Here $X=\vartheta_{1}$, $Y=\vartheta_{2}$, $Z=\vartheta_{3}$, $U=\vartheta_{0}$ are of degree $1$ and $W=\vartheta_{4}$ is of degree 0.
@@
fulfilling a compatibility condition [GHS, (4.8)]. Here $\operatorname{PL}(B)$ is the monoid of continuous maps $B\to\mathbb{R}$ that are piecewise affine with integral slopes on each cell of the polyhedral decomposition $\mathscr{P}$ of $B$. Explicitly, denoting by $\Sigma(\mathcal{Y})[1]$ the set of rays of $\Sigma(\mathcal{Y})$, the dual $\operatorname{PL}(B)^{*}$ is the quotient of $\mathbb{Z}^{\Sigma(\mathcal{Y})[1]}$ by the subspace of tuples pairing trivially with all piecewise linear functions. In our case of $Q=P=\mathbb{N}$ we take $\delta_{Q}=\operatorname{id}$ and $\delta_{B}$ the evaluation at the PL-function $\delta$ from Lemma 3.2 that vanishes on all bounded cells of the polyhedral decomposition $\mathscr{P}$ of $B$ and has slope $1$ on the unbounded rays.
It remains to check that the wall functions from (3.7) are homogeneous of degree 0 for the degree of monomials defined from $\delta_{Q}$ and $\delta_{B}$. Each wall function is a function in a monomial $t^{A}z^{-u_{\tau}}$ for a curve class $A$ and outgoing contact order $u_{\tau}$ for a type $\tau$ of a $1$-punctured map. In the notation of [GHS], this monomial has exponent $m=(\overline{m},m_{Q})$ with $\overline{m}=-u_{\tau}$, $m_{Q}=A\in P=\mathbb{N}$, and degree
@@
We have $\delta_{Q}(m_{Q})=A\in\mathbb{N}$, the intersection number $C]\cdot\mathcal{E}$ for the class of a curve $C$ contributing to the wall. In the second summand, $\nabla_{\overline{m}}\in\operatorname{PL}(B)^{*}$ is the directional derivative along $\overline{m}=-u_{\tau}$, so $\delta_{B}(\nabla_{\overline{m}})=\nabla_{\overline{m}}(\varphi)$. The balancing condition [ACGS, Cor. 2.30] applied with $s$ the section of $\overline{\mathcal{M}}_{X}^{\operatorname{gp}}$ defined by $\mathcal{E}$ then shows